What must be the distance between point charge q1 = 20μC and point charge q2 = −40μCfor the electrostatic force between them to have a magnitude of 10N?


1
Expert's answer
2021-04-19T17:06:08-0400

The distance can be found from Coulomb's law "F = k \frac{q_1 q_2}{r^2}".

From the last formula, "r = \sqrt{\frac{k q_1 q_2}{F}} = \sqrt{\frac{9 \cdot 10^9 \frac{N m^2}{C^2} \cdot 20 \cdot 10^{-6} C \cdot 40 \cdot 10^{-6} C}{10 N}} \approx 0.85 m"